# Learning to Play Two-Player Perfect-Information Games without Knowledge

### v4

@@ Abstract
In this paper, several techniques for learning game state evaluation functions by reinforcement are proposed. The first is a generalization of tree bootstrapping (tree learning): it is adapted to the context of reinforcement learning without knowledge based on non-linear functions. With this technique, no information is lost during the reinforcement learning process. The second is a modification of minimax with unbounded depth extending the best sequences of actions to the terminal states. This modified search is intended to be used during the learning process. The third is to replace the classic gain of a game ($+1$ / $-1$) with a reinforcement heuristic. We study particular reinforcement heuristics such as: quick wins and slow defeats ; scoring ; mobility or presence. The fourth is a new action selection distribution. The conducted experiments suggest that these techniques improve the level of play. We apply these different techniques to design program-players to the game of Hex (size $11$ and $13$) surpassing the level of Mohex 3HNN with reinforcement learning from self-play without knowledge.
keywords:
Machine Learning, Tree Search, Games, Game Tree Learning, Minimax, Heuristic, Action Distribution, Unbound Minimax, Hex.
@@ 1 Introduction
One of the most difficult tasks in artificial intelligence is the sequential decision making problem [1], whose applications include robotics and games. As for games, the successes are numerous. Machine surpasses man for several games, such as backgammon, checkers, chess, and go [2]. A major class of games is the set of two-player games in which players play in turn, without any chance or hidden information. This class is sometimes called two-player perfect information [^1] With some definitions, perfect information games include games with chance. It depends on whether one includes information about the future in perfect knowledge. games Mycielski [3] or also two-player combinatorial games. There are still many challenges for these games. For example, for the game of Hex, computers have only been able to beat strong humans since 2020 [4]. For general game playing [5] (even restricted to games with perfect information): man is always superior to machine on an unknown complex game (when man and machine have a relatively short learning time to master the rules of the game). In this article, we focus on two-player zero-sum games with perfect information, although most of the contributions in this article should be applicable or easily adaptable to a more general framework [^2] All the proposed techniques are directly applicable to the framework of perfect information single-agent sequential decision-making problems and in particular to solitaire games with perfect information. Tree learning, ordinal distribution, and reinforcement heuristics are applicable in the generalized cases (in case of $m$ agents, reinforcement heuristics can be valued in real $m$-space). Minimax variants are applicable in multi-player perfect-information games as variants of the paranoid algorithm [6]. Descent can be applied to learn value functions for stochastic perfect-information games by applying it to a large set of determinist games corresponding to the stochastic game where the chance was fixed and known before the start of the game. To use the corresponding learned heuristic, either a minimax at depth $1$ or an algorithm such as expectiminimax [7] must be used. However, for stochastic games and multiplayer games generalizations of the minimax variants should give better results in general..
@@
The first approaches used to design a game-playing program are based on a game tree search algorithm, such as minimax, combined with a handcrafted game state evaluation function based on expert knowledge. A notable use of this technique is the Deep Blue chess program [8]. However, the success of Deep Blue is largely due to the raw power of the computer, which could analyze two hundred million game states per second. In addition, this approach is limited by having to design an evaluation function manually (at least partially). This design is a very complex task, which must, in addition, be carried out for each different game. Several works have thus focused on the automatic learning of evaluation functions [9]. One of the first successes of learning evaluation functions is on the Backgammon game [10]. However, for many games, such as Hex or Go, minimax-based approaches, with or without machine learning, have failed to overcome human. Two causes have been identified [11]. Firstly, the very large number of possible actions at each game state prevents an exhaustive search at a significant depth (the game can only be anticipated a few turns in advance). Secondly, for these games, no sufficiently powerful evaluation function could be identified. An alternative approach to solve these two problems has been proposed, giving notably good results to Hex and Go, called Monte Carlo Tree Search and denoted MCTS [12, 13]. This algorithm explores the game tree non-uniformly, which is a solution to the problem of the very large number of actions. In addition, it evaluates the game states from victory statistics of a large number of random end-game simulations. It does not need an evaluation function. This was not enough, however, to go beyond the level of human players. Several variants of Monte Carlo tree search were then proposed, using in particular knowledge to guide the exploration of the game tree and/or random end-game simulations [13]. Recent improvements in Monte Carlo tree research have focused on the automatic learning of MCTS knowledge and their uses. This knowledge was first generated by supervised learning [14, 15, 16, 17, 18] then by supervised learning followed by reinforcement learning [19], and finally by only reinforcement learning [2, 20, 21]. This allowed programs to reach and surpass the level of world champion at the game of Go with the latest versions of the program AlphaGo [19, 2]. In particular, AlphaGo zero [2], which only uses reinforcement learning, did not need any knowledge to reach its level of play. This last success, however, required 29 million games of self-play (with 1,600 state evaluations per move). This approach has also been applied to chess [22]. The resulting program broke the best chess program (which is based on minimax). AlphaZero was subsequently reimplemented in Polygames [4], an open-source and multi-game program, which was notably applied to Hex and Havannah. In particular, at the 2020 Computer Olympiad, the gold medals for these games were won by Polygames.
@@
In this article [^3] Note that this paper is an extended, improved, and english version of [23]., we therefore focus on reinforcement learning within the minimax framework. We propose and asses new techniques for reinforcement learning of evaluation functions. Then, we apply them to design new program-players to the game of Hex (without using other knowledge than the rules of the game). We compare this program-player to Mohex 3HNN [16], the best Hex program, champion at Hex (size 11 and 13) of the $2018$ Computer Olympiad [24].
In the next section, we briefly present the game algorithms and in particular minimax with unbounded depth on which we base several of our experiments. We also present reinforcement learning in games, the game of Hex, the state of the art of game programs on this game, as well as other games on which experiments are performed. In the following sections, we propose different techniques aimed at improving learning performances and we expose the experiments carried out using these techniques. In particular, in Section 3, we extends the tree bootstrapping (tree learning) technique to the context of reinforcement learning without knowledge based on non-linear functions. In Section 4, we present a new search algorithm, a variant of unbounded minimax called descent, intended to be used during the learning process. In Section 5, we introduce reinforcement heuristics. Their usage is a simple way to use general or dedicated knowledge in reinforcement learning processes. We study several reinforcement heuristics in the context of different games. In Section 7, we introduce a new action selection distribution and we apply it with all the previous techniques to design program-players to the game of Hex (size 11 and 13) and compare them to Mohex 3HNN. Finally, in the last section, we conclude and expose the different research perspectives.
Important note: This article will soon receive an update to correct some minor issues and make main improvements (until then, contact me if you have any problems with this version).
@@ 2 Background and Related Work
In this section, we briefly present game tree search algorithms, reinforcement learning in the context of games and their applications to Hex and Chess (for more details about game algorithms, see [25]).
Games can be represented by their game tree (a node corresponds to a game state and the children of a node are the states that can be reached by an action). From this representation, we can determine the action to play using a game tree search algorithm. In order to win, each player tries to maximize his score (i.e. the value of the game state for this player at the end of the game). As we place ourselves in the context of two-player zero-sum games, to maximize the score of a player is to minimize the score of his opponent (the score of a player is the negation of the score of his opponent).
@@ 2.1 Game Tree Search Algorithms
The central algorithm is minimax which recursively determines the value of a node from the value of its children and the functions $\min$ and $\max$, up to a limit recursion depth. With this algorithm, the game tree is uniformly explored. A better implementation of minimax uses alpha-beta pruning [26, 25] which makes it possible not to explore the sections of the game tree which are less interesting given the values of the nodes already met and the properties of $\min$ and $\max$. Many variants and improvements of minimax have been proposed [27]. For instance, iterative deepening [28, 29] allows one to use minimax with a time limit. It sequentially performs increasing depth alpha-beta searches as long as there is time. It is generally combined with the move ordering technique [30], which consists of extending the best move from the previous search first, which accelerates the new search. Some variants perform a search with unbounded depth (that is, the depth of their search is not fixed) [31, 32, 33]. Unlike minimax with or without alpha-beta pruning, the exploration of these algorithms is non-uniform. One of these algorithms is the best-first minimax search [34]. To avoid any confusion with some best-first approaches at fixed depth, we call this algorithm Unbound Best-First Minimax, or more succinctly $\mathrm{UBFM}$. $\mathrm{UBFM}$ iteratively extends the game tree by adding the children of one of the leaves of the game tree having the same value as that of the root (minimax value). These leaves are the states obtained after having played one of the best sequences of possible actions given the current partial knowledge of the game tree. Thus, this algorithm iteratively extends the a priori best sequences of actions. These best sequences usually change at each extension. Thus, the game tree is non-uniformly explored by focusing on the a priori most interesting actions without exploring just one sequence of actions. In this article, we use the anytime version of $\mathrm{UBFM}$ [34], i.e. we leave a fixed search time for $\mathrm{UBFM}$ to decide the action to play. We also use transposition tables [35, 27] with $\mathrm{UBFM}$, which makes it possible not to explicitly build the game tree and to merge the nodes corresponding to the same state. Algorithm 1 is the used implementation of $\mathrm{UBFM}$ in this paper [^4] This implementation is a slight variant of Korf and Chickering algorithm. Their algorithm is very slightly more efficient but it offers less freedom: our algorithm behaves slightly differently depending on how we decide between two states having the same value. The exploration of states is identical between their algorithm and ours when with our algorithm equality is decided in deepest first. Our variant has been discovered independently of Korf and Chickering work..
@@
$T$ transposition table (contains functions as $v$ and/or $v^{\prime}$ ; depends on the used search algorithm)
@@
$v(s)$ value of state $s$ in the game tree according to the last tree search
$v^{\prime}(s,a)$ value obtained after playing action $a$ in state $s$
@@
                        $v^{\prime}(s,a)\leftarrow f\left(a(s)\right)$
@@
                  $v^{\prime}(s,a_{b})\leftarrow$ UBFM_iteration($a_{b}(s)$, $S$, $T$)
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v^{\prime}\left(s,a\right)}$
@@
             return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v^{\prime}\left(s,a\right)}$
@@
Algorithm 1 $\mathrm{UBFM}$ (Unbounded Best-First Minimax) algorithm : it computes the best action to play in the generated non-uniform partial game tree (see Table 1 for the definitions of symbols ; at any time $T=\left\{v^{\prime}(s,a)\,|\,s\in S,a\in\mathrm{actions}(s)\right\}$).
@@ 2.2 Learning of Evaluation Functions
Reinforcement learning of evaluation functions can be done by different techniques [9, 2, 20, 36]. The general idea of reinforcement learning of state evaluation functions is to use a game tree search algorithm and an adaptive evaluation function $f_{\theta}$, of parameter $\theta$, (for example a neural network) to play a sequence of games (for example against oneself, which is the case in this article). Each game will generate pairs $(s,v)$ where $s$ is a state and $v$ the value of $s$ calculated by the chosen search algorithm using the evaluation function $f_{\theta}$. The states generated during one game can be the states of the sequence of states of the game [10, 37]. For example, in the case of root bootstrapping (technique that we call root learning in this article), the set of pairs used during the learning phase is $D=\left\{\left(s,v\right)\,|\,s\in\mathbf{R}\right\}$ with $\mathbf{R}$ the set of states of the sequence of the game. In the case of the tree bootstrapping (tree learning) technique [37], the generated states are the states of the game tree built to decide which actions to play (which includes the states of the sequence of states of the game): $D=\left\{\left(s,v\right)\,|\,s\in\mathbf{T}\right\}$ with $\mathbf{T}$ the set of states of the partial game tree of the game. Thus, contrary to root bootstrapping, tree bootstrapping does not discard most of the information used to decide the actions to play. The values of the generated states can be their minimax values in the partial game tree built to decide which actions to play [37, 10]. Work on tree bootstrapping has been limited to reinforcement learning of linear functions of state features. It has not been formulated or studied in the context of reinforcement learning without knowledge and based on non-linear functions. Note that, in the case of AlphaGo Zero, the value of each generated state, the states of the sequence of the game, is the value of the terminal state of the game [2]. We call this technique terminal learning.
Generally between two consecutive games (between match phases), a learning phase occurs, using the pairs of the last game. Each learning phase consists in modifying $f_{\theta}$ so that for all pairs $(s,v)\in D$, $f_{\theta}(s)$ sufficiently approaches $v$ to constitute a good approximation. Note that, in the context of a variant, learning phases can use the pairs of several games. This technique is called experience replay [38]. Note that, adaptive evaluation functions $f_{\theta}$ only serve to evaluate non-terminal states since we know the true value of terminal states.
@@
There is another reinforcement learning technique for games: the Temporal Differences differences $\mathrm{TD}(\lambda)$ [10]. We can see the temporal differences as a kind of interpolation between root learning of a depth $1$ minimax and the terminal learning. A variant has been proposed [39], called $\mathrm{TD_{Leaf}}(\lambda)$, where the temporal difference technique is applied not to learn the value of the root but the value of the leaf of the principal variation of a minimax search (at any depth). A comparison between these techniques was made [37]. Finally, [40] describe a method for automatically tuning search-extension parameters, to decide which branches of the game tree must be explored during the search.

### v5

@@ Abstract
In this paper, several techniques for learning game state evaluation functions by reinforcement are proposed. The first is a generalization of tree bootstrapping (tree learning): it is adapted to the context of reinforcement learning without knowledge based on non-linear functions. With this technique, no information is lost during the reinforcement learning process. The second is a modification of Unbounded Best-First Minimax extending the best sequences of actions to the terminal states. This modified search is intended to be used during the learning process. The third is to replace the classic gain of a game ($+1$ / $-1$) with a reinforcement heuristic. We study particular reinforcement heuristics such as: quick wins and slow defeats ; scoring ; mobility or presence. The fourth is the completion technique which takes into account the resolution of states. The fifth is a new action selection distribution. The conducted experiments suggest that these new techniques improve the level of play.
Moreover, we combine all these techniques within an algorithm called Athénan. We compare Athénan with ExIt, a state-of-the-art algorithm for self-play reinforcement learning without knowledge. This comparison shows that Athénan is more performant. Finally, we apply Athénan to design program-players to the following games: Hex, Othello, Arimaa. The state-of-the-art level of theses games are surpassed without using prior knowledge. We also apply Athénan to the single-player game Morpion Solitaire which allowed to reach the state-of-the-art level without using prior knowledge too.
@@ 1 Introduction
One of the most difficult tasks in artificial intelligence is the sequential decision making problem [1], whose applications include robotics and games. As for games, the successes are numerous. Machine surpasses man for several games, such as backgammon, checkers, chess, and go [2]. A major class of games is the set of two-player games in which players play in turn, without any chance or hidden information. This class is sometimes called two-player perfect information [^1] With some definitions, perfect information games include games with chance. This is not the case in this article games [3] or also two-player combinatorial games. There are still many challenges for these games. For example, at the game of Hex, computers have only been able to beat strong humans since 2020 [4]. In the context of general game playing [5], i.e. the context of playing a unknown game with only a short time to learn how to play well: man is always superior to machine (even restricted to games with perfect information). In this article, we focus on two-player zero-sum games with perfect information. Note that the algorithms we propose also apply in the single-player case. We demonstrate this in the context of Morpion Solitaire.
@@
The first approaches used to design a game-playing program are based on a game tree search algorithm, such as minimax, combined with a handcrafted game state evaluation function based on expert knowledge. A notable use of minimax with an handcrafted evaluation is the Deep Blue chess program [6]. However, the success of Deep Blue is largely due to the raw power of the computer, which could analyze two hundred million game states per second. In addition, this approach is limited by having to design an evaluation function manually (at least partially). The design of evaluation functions is a very complex task, which must, in addition, be carried out for each different game. Several works have thus focused on automatic learning of evaluation functions [7]. One of the first successes about learning evaluation functions is on the Backgammon game [8]. However, for many games, such as Hex or Go, minimax-based approaches, with or without machine learning, have failed to overcome human. Two causes have been identified [9]. Firstly, the very large number of possible actions at each game state prevents an exhaustive search at a significant depth (the game can only be anticipated a few turns in advance). Secondly, for these games, no sufficiently powerful evaluation function could be identified. An alternative approach has been proposed to solve the two problems, called Monte Carlo tree search and denoted MCTS [10, 11]. MCTS explores the game tree non-uniformly, which is a solution to the problem of the very large number of actions. In addition, it evaluates the game states from victory statistics of a large number of random end-game simulations. Thus, it does not need an evaluation function. MCTS gived notably good results to Hex and Go. However, this was not enough to go beyond the level of human players. Several variants of Monte Carlo tree search were then proposed, using in particular knowledge to guide the exploration of the game tree and/or random end-game simulations [11]. Recent improvements in Monte Carlo tree search have focused on the automatic learning of MCTS knowledge and their uses. This knowledge was first generated by supervised learning [12, 13, 14, 15, 16] then by supervised learning followed by reinforcement learning [17], and finally by only reinforcement learning [2, 18, 19]. These learning-based improvements allowed programs to reach and surpass the level of world champion at the game of Go [17, 2]. In particular, the program AlphaGo Zero [2], which only uses reinforcement learning, did not need any knowledge to reach its level of play. This last success, however, required 29 million games of self-play (with 1,600 state evaluations per move). The AlphaGo Zero approach has also been generalized and applied to chess [20] under the name of AlphaZero. The resulting program broke the best chess program (based on minimax). Another state-of-the-art zero-knowledge reinforcement learning algorithm is the ExIt algorithm [18], also based on Monte Carlo tree search.
@@
In this article [^2] Note that this paper is an extended, improved, and english version of [21]., we therefore focus on reinforcement learning within the minimax framework. We propose and asses new techniques for reinforcement learning of evaluation functions, which we then combine together to form the Athénan algorithm. In particular, we show that Athénan makes it possible to go beyond the state-of-the-art of reinforcement learning without knowledge, in particular the ExIt algorithm, and the state-of-the-art dedicated to the following games: Hex, Arimaa, Othello, and Morpion Solitaire. Note that, we showed, in collaboration with Tristan Cazenave, in an another article produced after this article but published before, that Athénan is more efficient and performant than AlphaZero [22]. This further study complements the studies carried out within this article.
In the next section, we briefly present game algorithms and in particular Unbounded Best-First Minimax on which we base several of our experiments. We also present reinforcement learning in games and the games on which our experiments are performed. In the following sections, we propose different techniques aimed at improving learning performances and we expose the experiments carried out using these new techniques. In particular, in Section 3, we extends the tree bootstrapping (tree learning) technique to the context of reinforcement learning without knowledge based on non-linear functions. In Section 4, we present a new search algorithm, a variant of Unbounded Best-First Minimax called Descent, intended to be used during the learning process. In Section 6, we introduce reinforcement heuristics. Their usage is a simple way to use general or dedicated knowledge in reinforcement learning processes. We study several reinforcement heuristics in the context of different games. In Section 5, we introduce the completion technique for taking into account the resolution of states in the cases of Unbounded Minimax and Descent. Section 7 introduce a new action selection distribution.
Then, we combine previously introduced techniques into the algorithm called Athénan and apply it in different contexts. In Section 9, we apply Athénan to design program-players to the game of Hex (size $11$ and $13$) and compare them to Mohex 3HNN [14], the best Hex program (before Athénan), notably champion at Hex (size 11 and 13) during the $2018$ Computer Olympiad [23]. In Section 8, we compare Athénan with ExIt [18], the state-of-the-art reinforcement learning algorithm without knowledge at Hex. Then, in Section 10, we apply Athénan to surpass the dedicated state-of-the-art at Othello. In Section 11, we apply Athénan to surpass the dedicated state-of-the-art at Arimaa. In Section 17, we apply Athénan to reach the level of the dedicated state-of-the-art at Morpion Solitaire.
Moreover, in Section 13, we present the results of Athénan during the worldwide artificial intelligence competition on board games: the Computer Olympiad.
Finally, in Section 14, we conclude and expose the different research perspectives.
@@ 2 Background and Related Work
In this section, we briefly present game tree search algorithms, reinforcement learning in the context of games, and the games used in this paper (for more details about game algorithms, see [24]).
Games can be represented by their game tree (a node corresponds to a game state and the children of a node are the states that can be reached by an action). From this representation, we can determine the action to play using a game tree search algorithm. In order to win, each player tries to maximize his score (i.e. the value of the game state for this player at the end of the match). As we place ourselves in the context of two-player zero-sum games, to maximize the score of a player is to minimize the score of his opponent (the score of a player is the negation of the score of his opponent).
@@ 2.1 Game Tree Search Algorithms
The central algorithm is minimax which recursively determines the value of a node from the value of its children and the functions $\min$ and $\max$, up to a limit recursion depth. With minimax, the game tree is uniformly explored. A better implementation of minimax uses alpha-beta pruning [25, 24] which makes it possible not to explore the sections of the game tree which are less interesting given the values of the nodes already met and the properties of $\min$ and $\max$. Many variants and improvements of minimax have been proposed [26]. For instance, iterative deepening [27, 28] allows one to use minimax with a time limit. It sequentially performs increasing depth alpha-beta searches as long as there is time. It is generally combined with the move ordering technique [29], which consists of extending the best move from the previous search first, which accelerates the new search.
Some variants perform a search with unbounded depth (that is, the depth of their search is not fixed) [30, 31, 32]. Unlike minimax with or without alpha-beta pruning, the exploration of these algorithms is non-uniform. One of these algorithms is the best-first minimax search [33] [^3] At the time of the first experiments of this article, we were not aware that the Unbounded Best-First Minimax algorithm had already been proposed, so we rediscovered it independently during this work.. To avoid any confusion with some best-first approaches at fixed depth, we call this algorithm Unbound Best-First Minimax, or more succinctly $\mathrm{UBFM}$. $\mathrm{UBFM}$ iteratively extends the game tree by adding the children of one of the leaves of the game tree having the same value as that of the root (minimax value). These leaves are the states obtained after having played one of the best sequences of possible actions given the current partial knowledge of the game tree. Thus, that algorithm iteratively extends the a priori best sequences of actions. The best sequences usually change at each extension. Thus, the game tree is non-uniformly explored by focusing on the a priori most interesting actions without exploring just one sequence of actions.
In this article, we use the anytime version of $\mathrm{UBFM}$ [33], i.e. we leave a fixed search time for $\mathrm{UBFM}$ to decide the action to play. We also use transposition tables [34, 26] with $\mathrm{UBFM}$, which makes it possible not to explicitly build the game tree and to merge the nodes corresponding to the same state. (MCTS and Alpha-Beta in this paper also use transposition tables). Note that positive impact of transposition tables has been evaluated [35]. Algorithm 1 is an implementation of $\mathrm{UBFM}$ [^4] This implementation is a slight variant of Korf and Chickering algorithm. Both algorithms behave identically if the evaluation function gives a different value to each state and if transposition tables are not used. In practice, in the context of using transpositions, our variant improves win performance by a few percentage points [35]. If the evaluation function gives a different value to each state and transpositions are not used, Korf and Chickering algorithm is slightly faster (carelessly faster in some contexts).. $\mathrm{UBFM}$ has been improved within the algorithm called Unbounded Minimax with Safe Decision [36], denoted $\mathrm{UBFM}_{\mathrm{s}}$. This variant performs the same search in the game tree but plays a different action after performing this search. Instead of playing the action that leads to the state of best value, the action played is the one that was most selected from the current state during the search. This variant of Unbounded Minimax was originally proposed in the old, unpublished versions of this article [37]. It has been removed from this document to simplify it. The removed results about Unbounded Minimax with Safe Decision that appeared in the old versions of this article have been improved and detailed in the article that deals with its new introduction [36].
@@
$\mathrm{random}\left(\right)$ returns a uniformly random number from $[0,1]$
@@
$T$ transposition table (contains states labels as function $v$ ; depends on the used search algorithm)
@@
$v(s)$ value of state $s$ in the game tree
$v(s,a)$ value obtained after playing action $a$ in state $s$
@@
$c(s,a)$ completion value obtained after playing action $a$ in state $s$
@@
$r(s,a)$ resolution value obtained after playing action $a$ in state $s$
@@
                        $v(s,a)\leftarrow f\left(a(s)\right)$
@@
                  $v(s,a_{b})\leftarrow$ UBFM_iteration($a_{b}(s)$, $S$, $T$)
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v\left(s,a\right)}$
@@
             return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v\left(s,a\right)}$
@@
Algorithm 1 $\mathrm{UBFM}$ (Unbounded Best-First Minimax) algorithm : it computes the best action to play in the generated non-uniform partial game tree (see Table 1 for the definitions of symbols ; at any time $T=\left\{v(s,a)\,|\,s\in S,a\in\mathrm{actions}(s)\right\}$).
@@ 2.2 Learning of Evaluation Functions
Reinforcement learning of evaluation functions can be done by different techniques [7, 2, 18, 38]. The general idea of reinforcement learning of state evaluation functions is to use a game tree search algorithm and an adaptive evaluation function to play a sequence of matches (for example against oneself, which is the case in this article). Adaptive evaluation functions $f_{\theta}$ are usually neural networks, of parameter $\theta$. Each match generates pairs $(s,v)$ where $s$ is a game state and $v$ is the value of $s$ calculated by the chosen search algorithm using the evaluation function $f_{\theta}$. Such pairs are learned for updating $f_{\theta}$ in order to improve it. The set of states used for updating $f_{\theta}$ varies depending on the learning technique [8, 39]. For example, in the case of root bootstrapping (technique that we call root learning in this article), the set of pairs learned during the learning phase is $D=\left\{\left(s,v\right)\,|\,s\in\mathbf{R}\right\}$ with $\mathbf{R}$ the set of states of the match. In the case of the tree bootstrapping (tree learning) technique [39], the set of pairs learned during the learning phase is the set of states of the partial game tree built to decide which actions to play: $D=\left\{\left(s,v\right)\,|\,s\in\mathbf{T}\right\}$ with $\mathbf{T}$ the set of states of the search trees. Note that states of the match search trees include states of the match (since the states of the match are the roots of the search trees). Thus, contrary to root bootstrapping, tree bootstrapping does not discard most of the information used to decide the actions to play. The values of the generated states can be their minimax values in the partial game tree built to decide which actions to play [39, 8]. Work on tree bootstrapping has been limited to reinforcement learning of linear functions of state features. It has not been formulated or studied in the context of reinforcement learning without knowledge and based on non-linear functions. Note that, in the case of AlphaGo Zero, the learned states are the states of the sequence of the match (as with root learning), but the target value for each of these states is the value of the terminal state of the match [2]. We call that technique terminal learning.
Generally between two consecutive matches (between match phases), a learning phase occurs, using the pairs of the last match. Each learning phase consists in modifying $f_{\theta}$ so that for all pairs $(s,v)\in D$, $f_{\theta}(s)$ sufficiently approaches $v$ to constitute a good approximation. Note that, learning phases can also use the pairs of several matches. This technique is called experience replay [40]. Note that, adaptive evaluation functions $f_{\theta}$ only serve to evaluate non-terminal states since we know the true value of terminal states (it is important: using the exact terminal evaluation function improves confrontation performance [35]).
@@
There is another reinforcement learning technique for games: the Temporal Differences algorithm $\mathrm{TD}(\lambda)$ [8]. It corresponds to a root learning algorithm without search (i.e. with a search of depth $1$) where the target of a state is a weighted average of the values of its successor states in the match, parameterized by the constant $\lambda$. A variant of $\mathrm{TD}(\lambda)$ has also been proposed [41], called $\mathrm{TD_{Leaf}}(\lambda)$, where the temporal difference technique $\mathrm{TD}(\lambda)$ is modified by applying it with a minimax search (i.e. with a search of depth $d\geq 1$). A comparison between these techniques was made [39]. Finally, in [42] a method is described for automatically tuning search-extension parameters, to decide which branches of the game tree must be explored during the search.